# Learning Penalty for Optimal Partitioning via Automatic Feature Extraction

### v1

@@ Abstract
Changepoint detection identifies significant shifts in data sequences, making it important in areas like finance, genetics, and healthcare. The Optimal Partitioning algorithms efficiently detect these changes, using a penalty parameter to limit the changepoints number. Determining the appropriate value for this penalty can be challenging. Traditionally, this process involved manually extracting statistical features, such as sequence length or variance to make the prediction. This study proposes a novel approach that uses recurrent neural networks to learn this penalty directly from raw sequences by automatically extracting features. Experiments conducted on 20 benchmark genomic datasets show that this novel method surpasses traditional methods in partitioning accuracy in most cases.
@@
[…]
[xxx]organization=Northern Arizona University, city=Flagstaff, state=Arizona, country=United States
[…]
[yyy]organization=Université de Sherbrooke, city=Quebec, country=Canada
@@ 1 Introduction
Changepoint detection (also known as partitioning or segmentation) is crucial for identifying abrupt changes in data in various domains, including finance (Lattanzi and Leonelli, 2021), healthcare (Muggeo and Adelfio, 2010), network security (Tartakovsky et al., 2013), and environmental science (Reeves et al., 2007). This importance has driven the development of numerous methods over the past several decades. The literature review, gap in existing methods, and study contributions are based on the diagram in Figure 1.
@@
Figure 1: Literature review of the changepoint detection process in this study: (i) Selecting a detection method, (ii) Choosing a regularization method to control the changepoints number, (iii) Selecting a model type to predict the penalty, (iv) Set up the penalty prediction supervision, and (v) Choosing the model architecture
(i) (ii) Related Changepoint Detection Methods.
Changepoint detection methods can be divided into two categories: Deep Learning (DL) and Dynamic Programming (DP).
DL-based methods, which have evolved from earlier method Cumulative Sum (CUSUM) (Hinkley, 1971; James et al., 1987), including recent works (Li et al., 2024; Ermshaus et al., 2023) employ multilayer perceptrons (MLP) or k-nearest neighbors (KNN) as classification sliding windows to identify changepoints. However, these methods rely on local neighbors rather than the entire sequence and are overly complex for univariate data.
@@
In contrast, DP is well-suited for this context. These algorithms rely on a single hyperparameter to limit the changepoints number, which can be either: (a) a fixed changepoints number or (b) a penalty parameter to penalize the changepoint presence. Fixed changepoints number allow efficient DP algorithms (Auger and Lawrence, 1989; Bai and Perron, 2003; Killick et al., 2012; Rigaill, 2015) to locate them, but the changepoints number is rarely predetermined. More commonly used, penalty-based DP algorithms, such as Optimal Partitioning (OPART) (Jackson et al., 2005) and its variants—Pruned Exact Linear Time (PELT) (Killick et al., 2012), Functional Pruning Optimal Partitioning (FPOP) (Maidstone et al., 2016), and Labeled Optimal Partitioning (LOPART) (Hocking and Srivastava, 2023)—are widely regarded as the most effective. PELT and FPOP prune candidate changepoints efficiently, producing identical partitions as OPART, while LOPART extends OPART by incorporating predefined labels (the expected number of changepoints within specific location ranges) or defaulting to OPART operation when labels are absent.
@@
In OPART family algorithms, the penalty parameter plays a crucial role. A higher penalty imposes a stronger penalty on changepoint presence, leading to fewer detected changepoints, see Figure 2. While OPART use a fixed penalty, this study aims to predict this penalty value to enhance OPART accuracy. From here on, the OPART penalty will be referred to as $\lambda$. To gain a detailed understanding of how $\lambda$ works, please refer to the subsection “OPART Optimization Problem” in the appendix.
@@
Figure 2: Example demonstrating how different $\lambda$ values affect the OPART detection of changepoints. The labeled sequence contains four labels: three positive (one changepoint regions) and one negative (no changepoints region). Various $\lambda$ are tested in the OPART algorithm. Two types of label errors are considered: false positives (fp), where extra changepoints are detected in positive labels or any in negative labels, and false negatives (fn), where no changepoint is detected in a positive label.
(iii) (iv) Existing $\lambda$ prediction models limitation.
@@
Various models exist for predicting $\lambda$, including unsupervised (Schwarz, 1978; Akaike, 1974; Lavielle, 2005) and supervised such as linear (Rigaill et al., 2013) and trees (Drouin et al., 2017; Barnwal et al., 2022). But all of these models rely on a set of manually extracted statistical sequence features, yet technically the number of possible features is infinite, this process can result in the inclusion of many irrelevant features, and the potential omission of useful ones.
(iv) (v) Contribution.
@@
This study introduces a novel approach utilizing a recurrent neural network to automatically extract relevant features from sequences, which are then used to predict $\lambda$. Experimental results on benchmark datasets show that this approach surpasses previous methods in terms of OPART accuracy in the most cases.
@@
To ensure transparency and fairness, the code used to generate the results of this study is accessible at: https://github.com/lamtung16/ML_Changepoint_Detection_epigenomic_rnn.
@@ 2.1 Problem Setting
Each sequence is assigned labels, each indicating the expected changepoints count in a specific region, see the example in Figure 2 of a sequence having four labels. DL-based methods are not applicable here because they require supervision with exact changepoint locations.
The prediction models use interval regression.
The objective of this study is to predict the optimal $\lambda$ for each sequence to detect its changepoints. For each labeled sequence, the optimal $\lambda$ is not a single value, but rather exists within an interval $[\lambda_{l},\lambda_{u}]$. For example, in Figure 2, $\lambda=10$ is the optimal value, as it generates changepoints that align with expert labels; however, there exist $a,b\geq 0$ such that values of $\lambda$ within $[\lambda_{l},\lambda_{u}]=[10-a,10+b]$ could produce the same result. The algorithm for generating the optimal interval for each labeled sequence is detailed by Rigaill et al. (2013). This problem, where the prediction needs to fall within an interval, is called interval regression. In interval regression problem, there are 4 types of intervals: uncensored $(-\infty<\lambda_{l}=\lambda_{u}<\infty)$, interval-censored $(-\infty<\lambda_{l}<\lambda_{u}<\infty)$, left-censored $(-\infty=\lambda_{l}<\lambda_{u}<\infty)$ and right-censored $(-\infty<\lambda_{l}<\lambda_{u}=\infty)$.
@@ 2.2 Previous Work
This subsection focuses on the process of training supervised $\lambda$ prediction models. Before delving into the details, some noteworthy models are briefly highlighted, which, although not directly applicable to this study, are still relevant. Additionally, some unsupervised models, which have served as inspiration for supervised models, are mentioned.
@@
Adaptive Linear Penalty INference (ALPIN) (Truong et al., 2017) is a notable method for a special case of this supervision, where every single point in the sequence is a label, equivalent to label regions having length 1. However, since this study uses a more relaxed supervision with varying label region lengths, ALPIN is not applicable.
The Accelerated Failure Time (AFT) model (Wei, 1992) and its variants (Cai et al., 2009; Huang et al., 2006; Quinlan, 1986; Pölsterl et al., 2016), which are designed for censored outcomes, can be considered. Supervised changepoint detection setups involve interval-censored targets, whereas traditional AFT models are restricted to uncensored or right-censored intervals, making them not directly applicable. An important distinction is that in AFT models, the left-censored interval is defined as $(0=\lambda_{l}<\lambda_{u}<\infty)$, which is different from this study’s definition.
@@
The unsupervised models predicts $\lambda$ based on some sequence statistical features. Let $N$ denote the length and $\sigma$ the standard deviation of the sequence, with examples such as Baysian Information Criterion (BIC) (Schwarz, 1978) predicts $\lambda=\log N$ or $\lambda=\sigma^{2}\log N$. Akaike’s Information Criterion (AIC) (Akaike, 1974) predicts $\lambda=2p$, where $p$ represents a sequence feature (e.g., standard deviation, variance). Lavielle (2005), on the other hand, does not provide a closed-form formula but considers $\lambda$ to be dependent on both $N$ and $\sigma$.
The rest of this section will detail the training process of three supervised machine learning models: Maximum Margin Interval Regression (Rigaill et al., 2013), a linear model; Maximal Margin Interval Tree (MMIT) (Drouin et al., 2017); and AFT in XGBoost (Barnwal et al., 2022), which utilizes a tree-based framework (although AFT in XGBoost is an AFT model, it is applicable due to its ability to handle all types of non-negative intervals). These models divided the training process into 2 independent steps:
  • 1.
@@
  • 2.
    Step 2: Train a model $g(\cdot)$ takes into $\mathbf{x}$ to predict $\lambda$, $\lambda=g(\mathbf{x})$
@@ 2.2.1 Step 1: Feature Extraction
The manual feature extraction process, detailed in Rigaill et al. (2013), involves the following steps:
  • 1.
    Starting from the sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]$ with the mean $\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i}$, 2 additional vectors (residuals and differences) are generated:
@@
  • 3.
    Compute 8 statistics (sum, mean, standard deviation, and quantiles at 0%, 25%, 50%, 75%, 100%) from each transformed vector, yielding a feature vector of length 72.
  • 4.
@@
Undefined values, such as those caused by invalid operations (e.g., log of a negative number), is ignored.
@@ 2.2.2 Step 2: Model Training
After obtaining the feature vector $\mathbf{x}$ (up to 365 features), a model is selected and trained to predict $\lambda$.
@@
$\lambda=\mathcal{\bar{T}}(\mathbf{x})+\epsilon$
where $\mathcal{\bar{T}}(\mathbf{x})$ is the output from the decision tree ensemble based on the feature vector $\mathbf{x}$, and $\epsilon$ is a random error with a specified distribution (e.g., normal, log-normal, Weibull). To train the model, the error distribution is selected, and the AFT likelihood is constructed. Maximum likelihood estimation is then used for training. Optimal hyperparameters are chosen through cross-validation on the train set.
Figure 3: Example about the loss value of two loss functions is determined by comparing the prediction to the target. In the top plot, the Squared Error yields a value of 0 when the prediction $\hat{y}$ reaches the target point $y$. The bottom plot illustrates the Squared Hinge Error, where the loss value reaches 0 when the prediction $\hat{y}$ falls within the interval $[y_{l}+\epsilon,y_{u}-\epsilon]$
@@
This model, known as Maximum Margin Interval Regression, transforms the problem to predicting $\log\lambda$ instead of $\lambda$, as $\log\lambda$ can take any real value, unlike the non-negative $\lambda$. The model expresses the output as a linear combination of inputs:
$\log\lambda=\mathbf{x}\cdot\beta+\beta_{0}$
where $\beta$ is the parameter vector, $\mathbf{x}\cdot\beta$ is the dot product of $\mathbf{x}$ and $\beta$, and $\beta_{0}$ is the bias term. Its simplified version used in Hocking and Srivastava (2023) predicts $\log\lambda=\beta_{1}\times\log\log N+\beta_{0}$, where $N$ is the sequence length.
This model is trained by minimizing a specific loss function. Squared error (SE), commonly used in point estimate regression, is not suitable for this problem since each target is not a single point. Instead, the Hinge Error loss function, defined by Rigaill et al. (2013), generated from SE, is used. Unlike SE, which achieves 0 loss at a single point (when the prediction is identical with the target point), Hinge Error achieves 0 loss within an interval (when the prediction falls within the target interval), as shown in Figure 3. Using the ReLU function, the Hinge Error between prediction $\hat{y}$ and target interval $[y_{l},y_{u}]$ is expressed as:
$l(\hat{y},[y_{l},y_{u}])=\Big{(}\text{ReLU}(y_{l}-\hat{y}+\epsilon)\Big{)}^{p}+\Big{(}\text{ReLU}(\hat{y}-y_{u}+\epsilon)\Big{)}^{p}$ (1)
Here, the margin length $\epsilon$ is fixed ($\epsilon=1$ by default) and the loss type $p=2$ (Squared Hinge Error) are chosen.
@@
MMIT introduced by Drouin et al. (2017) – similar to regression trees by Breiman et al. (1984) – is applicable for $\lambda$ prediction as an interval regression model. Unlike Breiman et al. (1984), which minimizes Squared Error, MMIT minimizes the Hinge Error 1 within tree nodes to construct the prediction tree model:
$\log\lambda=\mathcal{T}(\mathbf{x})$
Here, $\mathcal{T}$ is the tree model. The optimal tree architecture (maximum depth, minimum sample split), loss type which is $p$ (either 1 or 2) in Hinge Error 1, and margin which is $\epsilon$ in Hinge Error 1, is selected through cross-validation.

### v2

@@ Abstract
Changepoint detection identifies significant shifts in data sequences, making it important in areas like finance, genetics, and healthcare. The Optimal Partitioning algorithms efficiently detect these changes, using a penalty parameter to limit the changepoints count. Determining the optimal value for this penalty can be challenging. Traditionally, this process involved manually extracting statistical features, such as sequence length or variance to make the prediction. This study proposes a novel approach that uses recurrent networks to learn this penalty directly from raw sequences by automatically extracting features. Experiments conducted on 20 benchmark genomic datasets show that this novel method generally outperforms traditional ones in changepoint detection accuracy.
@@ 1 Introduction
Changepoint detection (other words, partitioning or segmentation) is important for identifying abrupt changes in data in various domains, including finance [1], healthcare [2], network security [3], and environmental science [4]. This importance has driven the development of numerous methods over the past decades. The literature review, gap in existing methods, and study contributions are based on the diagram in Figure 1.
@@
Figure 1: Literature review of this study: (i) Selecting a detection method, (ii) Choosing a regularization method to control the changepoints number, (iii) Selecting a model type to predict the penalty, (iv) Set up the penalty prediction supervision, and (v) Choosing the model architecture.
Changepoint Detection Methods - Fig. 1 (i).
The methods can be divided into Deep Learning (DL) and Dynamic Programming (DP). Pure DL and DP approaches take different perspectives. DP methods treat the problem as an optimization task, aiming to find the optimal partition, whereas DL methods frame it as a binary classification problem, predicting whether each candidate point is a changepoint. So in terms of the underlying optimization, DP methods are optimal, while DL methods are heuristic. This category division approach is similar to that of Paul Fearnhead, the author of [5] (a DP method), and [6] (a DL method) where he classifies changepoint detection methods into traditional and neural network–based. DL-based methods, which inspired from a statistical method Cumulative Sum (CUSUM) [7, 8], including recent works [6, 9] employ multilayer perceptrons (MLP) or k-nearest neighbors (KNN) as classification sliding windows to identify changepoints. However, these methods rely on local neighbors rather than the entire sequence and are overly complex for univariate data. In contrast, DP is well-suited for this context.
Regularization - Fig. 1 (ii).
@@
The DP algorithms rely on a single hyperparameter to limit the changepoints number, which can be either: (a) a fixed changepoints number or (b) a penalty parameter to penalize the changepoint presence. Fixed changepoints number allow DP algorithms [10, 11, 5, 12] to locate them, but the changepoints number is rarely predetermined. More commonly used, penalty-based DP algorithms, such as Optimal Partitioning (OPART) [13] and its variants—Pruned Exact Linear Time (PELT) [5], Functional Pruning Optimal Partitioning (FPOP) [14], and Labeled Optimal Partitioning (LOPART) [15]—are widely regarded as the most effective. PELT and FPOP prune candidate changepoints efficiently, producing identical partitions as OPART, while LOPART extends OPART by incorporating predefined labels (the expected number of changepoints within specific location ranges) or defaulting to OPART when labels are absent. OPART makes no general assumptions about the time series itself — it is simply an optimization problem:
OPART Problem.
Given a sequence $\mathbf{d}\in\mathbb{R}^{N}$ and penalty parameter $\lambda\geq 0$, find the model parameter vector $\mathbf{m}\in\mathbb{R}^{N}$ that minimizes the following cost function:
$C(\textbf{d},\textbf{m})+\lambda\sum_{i=1}^{N-1}\mathbf{1}(m_{i}\neq m_{i+1}),$
where $\mathbf{1}(z)$ equals 1 when $z$ is True and 0 otherwise. The $C(\textbf{d},\textbf{m})$ formula depends on the type of change. For example, if the changes is the mean of normal distribution, then $C(\mathbf{d},\mathbf{m})=\sum_{i=1}^{N}(d_{i}-m_{i})^{2},$ or for changes in the variance, $C(\mathbf{d},\mathbf{m})=\sum_{i=1}^{N}\big((d_{i}-m_{i})^{2}-\sigma_{i}^{2}\big)^{2},$ where $\sigma_{i}^{2}$ is the variance of the segment corresponding to $m_{i}$.
@@
In OPART algorithms, the penalty plays a key role. A higher penalty imposes a stronger penalty on changepoint presence, leading to fewer changepoints, see Figure 2. This study aims to predict that penalty to improve the changepoint detection accuracy. From here on, that conceptual penalty will be referred to as $\lambda$.
@@
Figure 2: Example of how $\lambda$ affect the OPART changepoints. The sequence contains 4 labels: 3 positive (1 changepoint regions) and 1 negative (no changepoint region). Two types of errors are considered: false positives (fp), where extra changepoints are detected in any labels, and false negatives (fn), where no changepoint is detected in a positive label. In this examlpe, $\lambda=10$ is optimal.
Existing $\lambda$ prediction models - Fig. 1 (iii).
@@
Various models exist for predicting $\lambda$, including unsupervised [16, 17, 18] and supervised such as linear [19] and tree(s) [20, 21, 22]. But all of these models rely on a set of manually extracted sequence features, yet the number of possible features is infinite, this process can result in the inclusion of many irrelevant features, and the potential omission of useful ones.
Contribution - Fig. 1 (iv).
@@
This study proposes a novel method utilizing a recurrent network to automatically extract relevant features from sequences, which are then used to predict $\lambda$. Experiments on benchmark datasets show that the proposed method achieves higher partitioning accuracy than existing ones in most cases.
@@
To ensure transparency, the code used for this study is accessible at: https://github.com/lamtung16/ML_Changepoint_Detection_epigenomic_rnn.
@@ 2.1 Problem Setting
Each sequence is assigned multiple labels indicating the expected changepoints count in a specific region, see the example in Figure 2 of a sequence having four labels.
The $\lambda$ prediction models use interval regression.
This study aims to predict the optimal $\lambda$ for each sequence and then apply OPART with the predicted $\lambda$ to detect its changepoints. For each labeled sequence, the optimal $\lambda$ is not a single value, but rather is an interval $\lambda_{l},\lambda_{u}]$. For example, in Figure 2, $\lambda=10$ is the optimal value, as it generates changepoints that align with expert labels; however, there exist $a,b\geq 0$ such that values of $\lambda$ within [10a,10+b][10-a,10+b] could produce the same result, so [10a,10+b][10-a,10+b] is the target interval of that sequence. The algorithm for generating the target interval for each labeled sequence is detailed by Rigaill et al. [19]. This problem, where the model predicts the value that falls within the target interval, is known as interval regression. In interval regression, there are 4 types of target intervals $y_{l},y_{u}]$: uncensored $\infty<y_{l}=y_{u}<\infty]$, interval-censored $\infty<y_{l}<y_{u}<\infty]$, left-censored $\infty=y_{l}<y_{u}<\infty]$ and right-censored $\infty<y_{l}<y_{u}=\infty]$. In this study, each labeled sequence is associated with an interval $\lambda_{l},\lambda_{u}]$ where $0\leq\lambda_{l}\leq\lambda_{u}\leq\infty$.
@@ 2.2 Previous Work
This subsection focuses on the process of training supervised $\lambda$ prediction models. Before going into the details, some noteworthy models are highlighted, which, although not applicable to this study, are still relevant. Additionally, some unsupervised models, which have served as inspiration for supervised models, are mentioned.
@@
Adaptive Linear Penalty INference (ALPIN) [23] is a method for this supervision special case, where every point is a label, equivalent to label regions having length 1. However, since this study uses a more relaxed supervision with varying label region lengths, ALPIN is not applicable.
The Accelerated Failure Time (AFT) model [24] and its variants [25, 26, 27, 28], which are designed for censored outcomes, can be considered. Supervised changepoint detection setups involve interval-censored targets, whereas traditional AFT models are restricted to uncensored or right-censored intervals, making them not directly applicable.
@@
These models predicts $\lambda$ based on some sequence features. Let $\hat{\lambda}$ denote a prediction of $\lambda$. Let $N$ denote the sequence length, and $\sigma$ its standard deviation, with examples such as Baysian Information Criterion (BIC) [16] predicts $\hat{\lambda}=\sigma^{2}\log N$. Akaike’s Information Criterion (AIC) [17] predicts $\hat{\lambda}=2p$, where $p$ represents a sequence feature (e.g., standard deviation, variance). Lavielle [18] does not provide a closed-form formula but considers $\hat{\lambda}$ to be dependent on both $N$ and $\sigma$.
The rest of this section will detail the training process of three supervised machine learning models: Maximum Margin Interval Regression [19], a linear model; Maximal Margin Interval Tree (MMIT) [20]; and AFT in XGBoost [21], which utilizes a tree-based framework (although AFT in XGBoost is an AFT model, it is applicable due to its ability to handle all types of non-negative intervals). These models divided the training process into 2 sequential independent steps:
@@
  • Step 2: Train a model $g(\cdot)$ takes into $\mathbf{x}$ to predict $\hat{\lambda}=g(\mathbf{x})$
@@ 2.2.1 Step 1: Feature Extraction
The manual feature extraction process, detailed in [19], involves the following steps:
  • Starting from the sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]$ with the mean value $\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i}$, 2 additional sequences (residuals and differences) are generated:
@@
  • Compute 8 statistics (sum, mean, standard deviation, and quantiles at 0%, 25%, 50%, 75%, 100%) from each transformed vector, yielding a vector of length 72.
@@
The values caused by invalid operations (e.g., log of a negative number), is ignored.
@@ 2.2.2 Step 2: Model Training
After obtaining the feature vector $\mathbf{x}$ (up to 365 features), each instance $\{\mathbf{d},[\lambda_{l},\lambda_{u}]\}$ becomes $\{\mathbf{x},[\lambda_{l},\lambda_{u}]\}$, a selected model $g$ is trained to predict $\hat{\lambda}=g(\mathbf{x})$.
@@
$\hat{\lambda}=\mathbf{T}(\mathbf{x})+\varepsilon$
Here, $\mathbf{T}(\mathbf{x})$ denotes the ensemble of trees for feature vector $\mathbf{x}$, and $\varepsilon$ is a random error. The model is trained by minimizing the negative log-likelihood over all train instances:
$l_{AFT}\Big(\hat{\lambda},[\lambda_{l},\lambda_{u}]\Big)=-\log\Big[F_{Z}\big(\frac{\lambda_{u}-\hat{\lambda}}{\sigma}\big)-F_{Z}\big(\frac{\lambda_{l}-\hat{\lambda}}{\sigma}\big)\Big]$ (1)
where $\sigma>0$ is the scale parameter. The cumulative distribution function $F_{Z}(z)$ depends on the chosen distribution of $\varepsilon$:
  • normal: 00 $F_{Z}(z)=\frac{1}{2}\big(1+\text{erf}(\frac{z}{\sqrt{2}})\big)$
  • logistic: 00 $F_{Z}(z)=\frac{e^{z}}{1+e^{z}}$
  • extreme: 0 $F_{Z}(z)=1-e^{-\exp z}$
The model is trained using the gradient boosting algorithm [29] to minimize the loss value (1) over the training dataset with $S$ instances $\{\mathbf{x}^{i},[\lambda_{l}^{i},\lambda_{u}^{i}]\}_{i=1}^{S}$ given the distribution of $\varepsilon$ and the scale parameter $\sigma$:
$\sum_{i=1}^{S}l_{AFT}\Big(\hat{\lambda}^{i},[\lambda_{l}^{i},\lambda_{u}^{i}]\Big)=-\sum_{i=1}^{S}\log\Big[F_{Z}\big(\frac{\lambda_{u}^{i}-\mathbf{T}(\mathbf{x}^{i})}{\sigma}\big)-F_{Z}\big(\frac{\lambda_{l}^{i}-\mathbf{T}(\mathbf{x}^{i})}{\sigma}\big)\Big]$
@@
This model, known as Maximum Margin Interval Regression, transforms the problem to predicting $\log\hat{\lambda}$ instead of $\hat{\lambda}$, as $\log\hat{\lambda}$ can be any real value, unlike the non-negative $\hat{\lambda}$. The model expresses the output as a linear combination of inputs:
$\log\hat{\lambda}=\mathbf{x}\cdot\beta+\beta_{0}$ (2)
where $\beta$ is the parameter vector, $\mathbf{x}\cdot\beta$ is the dot product of $\mathbf{x}$ and $\beta$, and $\beta_{0}$ is the bias term. Its simplified version, used in [15], predicts $\log\hat{\lambda}=\beta_{1}\times\log\log N+\beta_{0}$, where $N$ is the sequence length.
This model is trained by minimizing a loss function, the Hinge Error, defined by Rigaill et al. [19], it achieves 0 loss when the prediction falls within the target interval. The error value between prediction $\log\hat{\lambda}$ and target interval $\log\lambda_{l},\log\lambda_{u}]$ is:
$l\Big(\log\hat{\lambda},[\log\lambda_{l},\log\lambda_{u}]\Big)=\begin{cases}(\log\lambda_{l}-\log\hat{\lambda}+\epsilon)^{2}&\mbox{if }\log\hat{\lambda}<\log\lambda_{l}+\epsilon,\\ 0&\mbox{if }\log\lambda_{l}+\epsilon\leq\log\hat{\lambda}\leq\log\lambda_{u}-\epsilon,\\ (\log\hat{\lambda}-\log\lambda_{u}+\epsilon)^{2}&\mbox{if }\log\hat{\lambda}>\log\lambda_{u}-\epsilon.\end{cases}$
where $\epsilon\geq 0$ is the margin length ($\epsilon=0$ by default). The Hinge Error becomes Square Error when $\lambda_{l}=\lambda_{u}$ and $\epsilon=0$ (or $\epsilon=\frac{\log\lambda_{l}+\log\lambda_{u}}{2}$), or in other word, target interval becomes target point. Drouin et al. [20] proposed a similar loss function:
$l\Big(\log\hat{\lambda},[\log\lambda_{l},\log\lambda_{u}]\Big)=\begin{cases}|\log\lambda_{l}-\log\hat{\lambda}+\epsilon|&\mbox{if }\log\hat{\lambda}<\log\lambda_{l}+\epsilon,\\ 0&\mbox{if }\log\lambda_{l}+\epsilon\leq\log\hat{\lambda}\leq\log\lambda_{u}-\epsilon,\\ |\log\hat{\lambda}-\log\lambda_{u}+\epsilon|&\mbox{if }\log\hat{\lambda}>\log\lambda_{u}-\epsilon.\end{cases}$
Using the ReLU function, the general loss function can be written in this short form:
$l(\log\hat{\lambda},[\log\lambda_{l},\log\lambda_{u}])=\Big(\text{ReLU}(\log\lambda_{l}-\log\hat{\lambda}+\epsilon)\Big)^{p}+\Big(\text{ReLU}(\log\hat{\lambda}-\log\lambda_{u}+\epsilon)\Big)^{p},$ (3)
where $\epsilon\geq 0$ is the margin length and $p\in\{1,2\}$ is the loss type.
To estimate the parameters $\beta$ and $\beta_{0}$ in (2), the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) [30] with respect to $\beta$ and $\beta_{0}$ is employed to minimize the Hinge Error (3) over the train dataset having $S$ instances $\{\mathbf{x}^{i},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\}_{i=1}^{S}$:
$\sum_{i=1}^{S}l\Big(\log\hat{\lambda}^{i},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\Big)=\sum_{i=1}^{S}l\Big(\mathbf{x^{i}}\cdot\beta+\beta_{0},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\Big)$
@@
MMIT introduced by Drouin et al. [20] is applicable for $\lambda$ prediction as an interval regression model. The MMIT model is:
$\log\hat{\lambda}=\mathcal{T}(\mathbf{x})$
The tree $\mathcal{T}$ architecture follows the same structure as CART (Classification And Regression Tree) [31]. The only difference lies in the regression value for each leaf which is a set of $Q$ instances $\{\mathbf{x}^{i},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\}_{i=1}^{Q}$: instead of taking the mean of all targets like CART does, MMIT chooses a constant value $c$ that minimizes the mean Hinge Error (3) between this constant and the target intervals:
$\operatorname*{argmin}_{c}\frac{1}{Q}\sum_{i=1}^{Q}l\big(c,[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\big),\quad c\in\mathbf{R}$
A heuristic algorithm is used to speed up the estimation of the mean value for each leaf by considering a finite set of candidates, rather than solving for the optimal $c$:
$\operatorname*{argmin}_{c}\frac{1}{Q}\sum_{i=1}^{Q}l\big(c,[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\big),$
$\text{where }c\in\{\log\lambda_{l}^{i},\log\lambda_{u}^{i},\frac{\log\lambda_{l}^{i}+\log\lambda_{u}^{i}}{2}|\log\lambda_{l}^{i}>-\infty,\log\lambda_{u}^{i}<\infty\}_{i=1}^{Q}$
The training procedure for determining the optimal tree architecture follows the same operations as in CART (binary optimization) to minimize the Hinge Error (3) over the train dataset having $S$ instances $\{\mathbf{x}^{i},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\}_{i=1}^{S}$:
$\sum_{i=1}^{S}l\Big(\log\hat{\lambda}^{i},[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\Big)=\sum_{i=1}^{S}l\Big(\mathcal{T}(\mathbf{x}^{i}),[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]\Big)$
Recurrent core: $h_{t}=f(d_{[t-k+1:t]};h_{t-1})\in\mathbb{R}^{m}$
sequence feature $k$ $m$ $h_{k-1}$ $h_{t}$ used in
length $k\in\mathbf{N^{+}}$ 1 k1k-1 $h_{t-1}+1$ [16, 17, 18, 20, 21, 19, 15, 32]
mean 1 1 0 $h_{t-1}+\frac{d_{t}-h_{t-1}}{t}$ [20, 21, 19]
min 1 1 $\infty$ $\min(d_{t},h_{t-1})$ [20, 21, 19]
max 1 1 -$\infty$ $\max(d_{t},h_{t-1})$ [20, 21, 19]
sum 1 1 0 $h_{t-1}+d_{t}$ [20, 21, 19]
cumulative absolute 1 1 0 $h_{t-1}+|d_{t}|$ [20, 21, 19]
cumulative squares 1 1 0 $h_{t-1}+d_{t}^{2}$ [20, 21, 19]
difference sum 2 1 0 $|d_{t}-d_{t-1}|+h_{t-1}$ [32, 20, 21, 19]
value range 1 2 $\infty$ $\min(d_{t},h_{t-1}^{(1)})$ [32, 20, 21, 19]
-$\infty$ $\max(d_{t},h_{t-1}^{(2)})$
Table 1: Formula of recurrent core for each feature approximation.